# Uniform approximation of Betti numbers

### v1

@@ Introduction
Since Grothendieck introduced étale cohomology, proving that the Betti numbers of $\ell$-adic cohomology of algebraic varieties over a separably closed field are independent of $\ell$ remains a deep open problem in positive characteristic. The case of smooth and proper algeraic varieties is reached by weight techniques, using Deligne’s proof of the Weil conjectures, but there is not any other general result in this direction. The next best thing consists in providing uniform bounds of Betti numbers. This can be done by dévissage from the case of smooth and proper algebraic varieties, and then applying descent techniques together with de Jong’s alteration theorem, as explained by Katz [Kat01, Theorem 5] who credits the result to de Jong and Berthelot. There are results for non-constant coefficients as well, working with compatible families.
The present notes give new proofs of existence of such uniform bound, using motivic techniques. We apply Lefschetz’s principle of approximating a quasi-projective scheme $X$ by a suitable hyperplane section. Of course, this needs adjustments in order to take into account the singularities of $X$ as well as the fact that we want the choice of the hyperplane section to be independent of the cohomology we work with (e.g. $\ell$-adic cohomology of schemes of finite type over a separably closed field $k$, with $\ell$ distinct from the characteristic of $k$).
To be more precise, the general strategy goes as follows. In this introduction, in order to simplify things, we work over an infinite field $k$ of characteristic $p$. Using Jouanolou’s trick [Jou73, Lemme 1.5], since the formation of Voevodsky motives is $\mathbf{A}^{1}$-homotopy invariant, one can focus on affine schemes of finite type over $k$. For such a scheme $X$ of dimension $d$, one defines the notion of Leschetz subscheme: a closed subscheme $Z$ of $X$ of dimension d1d-1, with affine complement, that contains all irreducible components of $X$ of dimension <d<d and the singular locus of $X$, as well as a hyperplane section satisfying a suitable transversality condition, expressed in the language of the six operations for motivic sheaves; see Definition 3.2 below. We define the motive of the pair $(X,Z)$ through the cofiber sequence
@@
in Voevodsky’s category $\mathit{DM}(\mathrm{Spec}(k))$ (of étale motives). We prove that, for any prime number $\ell\neq p$, the $\ell$-adic realization of $M(X,Z)$ is concentrated in degree d-d and that its unique cohomology group is a free $\mathbf{Z}_{\ell}$-module of finite type, using the theory of $\ell$-adic perverse sheaves [^1] This is where the fact that we work with affine schemes matters: the main tool is then Artin’s vanishing theorem.; see Lemma 3.4. Since the $\ell$-adic realization of motivic sheaves is symmetric monoidal, it is compatible with the formation and the computation of traces of endomorphisms of constructible motives, in particular of Euler characteristics (these are traces of identities). In this case, this means that the Euler characteristic of M(X,Z)[d]M(X,Z)[-d], seen as a motive, coincides with the rank $\mathit{rk}(X,Z)$ of the unique non-trivial cohomology group of its $\ell$-adic realization, proving that the latter is independent of $\ell$. This yields isomorphisms
@@
This means that the $i$th $\ell$-adic Betti number of the affine scheme $X$ is bounded by $\mathit{rk}(X_{i},X_{i-1})$ for all $\ell\neq p$; see Theorem 3.6.
However, our proof does more than providing a bound that is independent of $\ell$: it provides a uniform bound that is valid in any realization of Voevodsky’s motivic sheaves that has a theory of perverse sheaves (including the properties that Verdier duality induces an equivalence on constructible objects and that Artin’s vanishing theorem holds); see Definition 2.7. By developing this axiomatic approach, we also try to reach a wide enough level of generality, geometrically: instead of working over a field, we develop enough material to prove all the main ingredients over a scheme of finite type over a field: what is true over an infinite field is in fact true generically, possibly after pulling back along a flat finite cover of a dense open subset of any algebraic variety (from this point of view, we can work over any field $k$). Because what is at stake, after all, is the problem of finding big enough families of motivic sheaves that are known to be realized into ordinary sheaves or into perverse sheaves. This is what we do at the end of this article, thus providing constructions of $t$-structures on $\mathit{DM}(X,\mathbf{Q})$ that are candidates for the perverse motivic $t$-structure. These constructions, by finding nice generators for categories of motivic sheaves, may be seen as a way to provide uniform bounds of Betti numbers of all sorts, not only for the cohomology of algebraic varieties, but for the cohomology of any constructible motivic sheaf (this includes cohomology with compact support or not, as well as nearby cycles, thanks to Ayoub’s work [Ayo07]), possibly in relative situations. Finding explicit generators for such $t$-structures is also part of the standard tools to develop the theory of Nori motives and to relate them with Voevodsky motives in characteristic zero. Therefore, the results of the present text open the way to a uniform theory of Nori motives [HMS17, Ara05, MI19], possibly over any field: instead of working with a single cohomology theory, we can now consider the family of all cohomology theories appearing in a six-functors formalism in which there is a good notion of perverse sheaves. This is strongly related with Ayoub’s constructions [Ayo23] of new Weil cohomologies in positive characteristic and their associated Hopf algebroids.
@@ 1. Motivic shadow of the weak Lefschetz Theorem
We consider a $\mathbf{Z}$-linear six-functors formalism $D$ in the sense of Voevodsky-Ayoub (see [DG22]), taking values in presentable $\infty$-categories, defined on the category of schemes locally of finite presentation over a base scheme $S$. We assume furthermore that $D$ has the property of étale descent: for any surjective étale map $f:X\to Y$, the pullback functor
$f^{*}=f^{!}\colon D(Y)\to D(X)$
is conservative – equivalently, the assignment $X\mapsto D(X)$ is a sheaf of $\infty$-categories with respect to the étale topology. We assume further more that, for any $X$, there is a full stable subcategory $D_{c}(X)$ of $D(X)$, whose objects are called constructible objects with the following properties:
@@
    for any $X$ and any $F$ in $D_{c}(X)$, there is a dense open subscheme $U\subset X$ such that $F_{|U}$ is dualizable (in the sense of [Lur17, Section 4.6.1]).
@@
    for any $X$, any dualizable object of $D(X)$ is constructible.
@@
Let $\mathit{DM}(X)=\mathrm{Ind}(\mathit{DM}_{h,lc}(X))$ the $\infty$-category of ind-objects on the $\infty$-category of locally constructible objects of the $\infty$-category of $h$-motives in the sense of [CD16, Def. 6.3.1]. Then $\mathit{DM}$ satisfies all the properties above, at least if we restrict to noetherian schemes (but this is not even necessary); this follows easily from see [CD16, Theorem 6.3.26]. For any commutative ring $\Lambda$, there is a $\Lambda$-linear version $\mathit{DM}(X,\Lambda)$ defined as the $\Lambda$-modules in $\mathit{DM}(X)$. There is a rigidity theorem for $\mathit{DM}$ in the following sense: if $\Lambda$ is a commutative ring of characteristic $n>0$ with $n$ prime to the residue characteristics of $X$, there is a canonical equivalence between compact $\Lambda$-modules in $\mathit{DM}(X)$ and ordinary constructible sheaves of $\Lambda$-modules of finite tor-dimension over the small étale site of $X$.
@@
is then a six-functors formalism that satisfies all properties listed in paragraph 1.1. We observe that the dualizable objects are then exactly those complexes whose cohomology objects are what Deligne calls “faisceaux locallement constants tordus”. Indeed, the functor
@@
Let $\mathcal{U}$ be an ultrafilter on the set of prime numbers. We define for any $X$:
@@
(where the colimit and the product are taken in the $\infty$-category of (stable) $\infty$-categories) and
@@
If the set of residue characteristics of $X$ is finite, $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})$ coincides with the $\mathcal{U}$-indexed ultraproduct of the ordinary categories of constructible sheaves $D^{b}_{c}(X,\mathbf{F}_{\ell})$. It the set of residue characteristics is finite (e.g. is X is of finite type over a semi-local scheme), then $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})$ coincides with the ultraproduct of the usual categories of étale sheaves $D(X,\mathbf{F}_{\ell})$, where $\ell$ runs over all primes distinct from the residue characteristics of $X$. The corresponding cohomology is then the ultraproduct in the usual sense of étale cohomology groups:
@@
that are compatible with pullbacks $f^{*}$ for all $f$ as well with direct image with compact support $f_{!}$ for any $f$ separated and of finite presentation. In particular, on noetherian schemes, such morphisms must preserve constructible objets (they preserve dualizable objects, which implies that they preserve constructible objects by noetherian induction). One can prove that $\mathit{DM}$ is the initial object among $\mathbf{Z}$-linear six-functors formalisms satisfying étale descent [^2] One can see that $\mathit{DM}_{h}$ is initial among $\mathbf{Z}$-linear six-functors formalisms satisfying étale hyperdescent: since we know what is the initial object by virtue of Drew and Gallauer’s [DG22, Theorem 7.14], we see easily that it suffices to prove it for $\mathbf{Q}$-linear coefficients, which follows right away from [CD19, Theorem 16.1.2 and 16.2.18], and for finite coefficients, which follows from the combination of rigidity theorems: the one recalled in Example 1.2 together with those stated in [Ayo14, Bac21]. Since for schemes of finite type over $\mathbf{Z}$, there is no difference between étale descent and étale hyperdescent, and since $\mathit{DM}_{h}(X)$ is then compactly generated (hence $\mathit{DM}_{h}(X)=\mathit{DM}(X)$), we conclude by noetherian approximation together with continuity [CD16, Theorem 6.3.12] that $\mathit{DM}$ is the initial object among $\mathbf{Z}$-linear six-functors formalisms satisfying étale descent.. In particular, for any other six-functor formalism $D$, there is a unique morphism
@@
We call an object $F$ of $D(X)$ strictly of geometric origin if there is a constructible motivic sheaf $M$ in $\mathit{DM}_{c}(X)$ such that $F\cong r_{D}(M)$ in $D(X)$ (in particular, such an $F$ is constructible). We will say that $D$ behaves well over $S$ if the functor $r_{D}$ commutes with functors of the form $f_{*}$ for any morphism of finite presentation $f$. This is true in the case where $D(X)=D(X,\mathbf{Z}_{\ell})$ for any $S$ (see [CD16, Theorem 7.2.11]), or $D(X)=D(X,\mathbf{F}_{\mathcal{U}})$ (exercise), or for any $D$ in the case where $S$ is of finite type over a field $k$, and $D$ is defined over schemes of finite type over $k$ (using de Jong’s resolution of singularities by alteration as well as the fact that $DM$ is a sheaf for the $h$-topology).
@@
The following statement, is a generalization of Deligne’s generic base change theorem for torsion étale sheaves [Del77, Th. Finitude, 1.9]. With this new degree of generality, it may be seen as an independence of $D$ (in particular, independence of $\ell$) result as well.
@@
The proof of the generic base change formula really uses étale descent: this implies that, for any quasi-finite morphism $f$. the functors $f_{!}$ and $f_{*}$ are conservative, which is a crucial ingredient in Deligne’s proof. It is not known whether the generic base change formula is true or not for more universal theories of motivic sheaves in which we only have Nisnevich descent, such as Voevodsky’s and Morel’s $\mathit{SH}$.
@@
If moreover $D$ behaves well over $S$ (e.g. if $S$ is of finite type over a field $k$ and $D$ is defined on schemes of finite type over $k$) and if $F$ is strictly of geometric origin, then the open subscheme $U\subset S$ in the statement of Theorem 1.11 may be chosen indepently of $D$.
@@
Let $M$ in $\mathit{DM}_{c}(X)$ such that $F\cong r_{D}(M)$ in $D(X)$. Apply the generic base change formula to $M$, which gives an open subscheme $U\subset S$ such that the conclusion of Theorem 1.11 holds in $\mathit{DM}$. Then, since $D$ behaves well over $S$, the conclusion of Theorem 1.11 holds for $F$ in $D$ with respect to the motivically chosen $U\subset S$: since $r_{D}\colon\mathit{DM}\to D$ commutes with the six operations, all the base change isomorphisms involving $M$ in $\mathit{DM}$ will yield analogous base change isomorphisms involving $F$ in $D$. ∎
Remark 1.14.
In fact, a close examination of the proof of the generic base change formula shows that, for a given constructible $F$ in $D(X)$ and a morphism of finite presentation $f:X\to Y$ over some base $S$, one can choose the open subset $U\subset S$ so that the conclusion of the generic base change theorem holds for $f_{*}(F)$ in $D$ so that, for any morphism of six-functors formalisms $\varphi\colon D\to D^{\prime}$ (with $D^{\prime}$ satisfying the assumption of 1.1 as well), the canonical map
$\varphi(f_{*}(F))\to f_{*}(\varphi(F))$
becomes an isomorphism over $U$ and the conclusion of the generic base change theorem holds for $\varphi(F)$ with respect to $f$ in $D^{\prime}$. This means that the hypothesis that $D$ is well behaved is not necessary in the previous corollary. This is also proves that all six-functors formalisms satisfying the assumptionms of 1.1 are well behaved for schemes of finite type over a field.
@@
with $F_{T}$ the pullback of $F$ along the projection $X_{T}\to X$. Moreover, if ever $F$ is strictly of geometric origin, this open subset $U$ can be chosen independently of $D$.
@@
in which the vertical map on the right hand side is an isomorphism by the computation we just did. Since the middle vertical map is the identity, this proves that the vertical map on the left hand side is an isomorphism. The fact the open subscheme $U\subset\mathbf{P}^{\wedge}$ may be chosen independently if ever the sheaf $F$ is strictly of geometric origin follows from Corollary 1.13. ∎

### v2

@@ Introduction
Since Grothendieck introduced étale cohomology, proving that the Betti numbers of $\ell$-adic cohomology of algebraic varieties over a separably closed field are independent of $\ell$ remains a deep open problem in positive characteristic. The case of smooth and proper algeraic varieties is reached by weight techniques, using Deligne’s proof of the Weil conjectures, but there is not any other general result in this direction. The next best thing consists in providing uniform bounds of Betti numbers. This can be done by dévissage from the case of smooth and proper algebraic varieties, and then applying descent techniques together with de Jong’s alteration theorem, as explained by Katz [Kat01, Theorem 5] who credits the result to de Jong and Berthelot. There are results for non-constant coefficients as well, working with compatible families [KL85, Kat01, Org19].
The present notes give new proofs of existence of such uniform bound, using motivic techniques (in particular, we do not only prove independence of $\ell$ results but rather independence against a large class of Weil cohomologies). We apply Lefschetz’s principle of approximating a quasi-projective scheme $X$ by a suitable hyperplane section. Of course, this needs adjustments in order to take into account the singularities of $X$ as well as the fact that we want the choice of the hyperplane section to be independent of the cohomology we work with (e.g. $\ell$-adic cohomology of schemes of finite type over a separably closed field $k$, with $\ell$ distinct from the characteristic of $k$, or Berthelot’s rigid cohomology if $k$ is perfect of characteristic $p>0$).
To be more precise, the general strategy goes as follows. In this introduction, in order to simplify things, we work over an infinite field $k$ of characteristic $p$. Using Jouanolou’s trick [Jou73, Lemme 1.5], since the formation of Voevodsky motives is $\mathbf{A}^{1}$-homotopy invariant, the motive of any quasi-projective scheme is isomorphic to the motive of an affine one. Therefore, we can focus on affine schemes of finite type over $k$. For such a scheme $X$ of dimension $d$, one defines the notion of Leschetz subscheme: a closed subscheme $Z$ of $X$ of dimension d1d-1, with affine complement, that contains all irreducible components of $X$ of dimension <d<d and the singular locus of $X$, as well as a hyperplane section satisfying a suitable transversality condition, expressed in the language of the six operations for motivic sheaves; see Definition 3.2 below. We define the motive of the pair $(X,Z)$ through the cofiber sequence
@@
in Voevodsky’s category $\mathit{DM}_{c}(\mathrm{Spec}(k))$ (of constructible étale motives). We prove that, for any prime number $\ell\neq p$, the $\ell$-adic realization of $M(X,Z)$ is concentrated in degree d-d and that its unique non-trivial cohomology group is a free $\mathbf{Z}_{\ell}$-module of finite type, using the theory of $\ell$-adic perverse sheaves [^1] This is where the fact that we work with affine schemes matters: the main tool is then Artin’s vanishing theorem in order to prove Beilinson’s transversality Theorem 2.22 below.; see Lemma 3.4. Since the $\ell$-adic realization of motivic sheaves is symmetric monoidal, it is compatible with the formation and the computation of traces of endomorphisms of constructible motives, in particular of Euler characteristics (these are traces of identities). In this case, this means that the Euler characteristic of M(X,Z)[d]M(X,Z)[-d], seen as a motive, coincides with the rank $\mathit{rk}(X,Z)$ of the unique non-trivial cohomology group of its $\ell$-adic realization, proving that the latter is independent of $\ell$. This yields isomorphisms
@@
This means that the $i$th $\ell$-adic Betti number of the affine scheme $X$ (both $\mathbf{Q}_{\ell}$- or with $\mathbf{F}_{\ell}$-coefficients) is bounded by $\mathit{rk}(X_{i},X_{i-1})$ for all $\ell\neq p$; see Theorem 3.6.
However, our proof does more than providing a bound that is independent of $\ell$: it provides a uniform bound that is valid in any realization of Voevodsky’s motivic sheaves that has a theory of perverse sheaves (including the properties that Verdier duality induces an equivalence on constructible objects and that Artin’s vanishing theorem holds); see Definition 2.7. By developing this axiomatic approach, we also try to reach a wide enough level of generality, geometrically: instead of working over a field, we develop enough material to prove all the main ingredients over a scheme of finite type over a field: what is true over an infinite field is in fact true generically, possibly after pulling back along a flat finite cover of a dense open subset of any algebraic variety (from this point of view, we can work over any field $k$). Because what is at stake, after all, is the problem of finding big enough families of motivic sheaves that are known to be realized into ordinary sheaves or into perverse sheaves. This is what we do at the end of this article (Prop. 3.9 and 3.17), thus providing constructions of $t$-structures on $\mathit{DM}(X,\mathbf{Q})$ that are candidates for the perverse motivic $t$-structure (3.18). These constructions, by finding nice generators for categories of motivic sheaves, may be seen as a way to provide uniform bounds of Betti numbers of all sorts, not only for the cohomology of algebraic varieties, but for the cohomology of any constructible motivic sheaf (this includes cohomology with compact support or not, as well as nearby cycles, thanks to Ayoub’s work [Ayo07]), possibly in relative situations. Finding explicit generators for such $t$-structures is also part of the standard tools to develop the theory of Nori motives and to relate them with Voevodsky motives in characteristic zero. Therefore, the results of the present text open the way to a uniform theory of Nori motives [HMS17, Ara05, IM19], possibly over any field: instead of working with a single cohomology theory, we can now consider the family of all cohomology theories appearing in a six-functors formalism in which there is a good notion of perverse sheaves. This is strongly related with Ayoub’s constructions [Ayo23] of new Weil cohomologies in positive characteristic and their associated Hopf algebroids.
@@ 1. Motivic shadow of the weak Lefschetz Theorem
We consider a $\mathbf{Z}$-linear six-functors formalism $D$ in the sense of We assume further more that, for any $X$, there is a full stable subcategory $D_{c}(X)$ of $D(X)$, whose objects are called constructible objects with the following properties:
@@
    for any $X$ and any $F$ in $D_{c}(X)$, there is a dense open subscheme $U\subset X$ such that $F_{|U}$ is dualizable (in the sense of [Lur17, Section 4.6.1]);
@@
    for any $X$, any dualizable object of $D(X)$ is constructible;
  • (iv)
    constructible objects form an hypersheaf of $\infty$-categories for the étale topology.
@@
Let $\mathit{DM}(X)=\mathrm{Ind}(\mathit{DM}_{h,lc}(X))$ the $\infty$-category of ind-objects on the $\infty$-category of locally constructible objects of the $\infty$-category of $h$-motives in the sense of [CD16, Def. 6.3.1]. Then $\mathit{DM}$ satisfies all the properties above, at least if we restrict to noetherian schemes (but this is not even necessary); this follows easily from see [CD16, Theorem 6.3.26]. For any commutative ring $\Lambda$, there is a $\Lambda$-linear version $\mathit{DM}(X,\Lambda)$ defined as the $\Lambda$-modules in $\mathit{DM}(X)$.
If $\Lambda$ is a commutative ring of characteristic $n>0$, there is a rigidity theorem in the following sense: for $n$ prime to the residue characteristics of $X$, there is a canonical equivalence between compact $\Lambda$-modules in $\mathit{DM}(X)$ and ordinary constructible sheaves of $\Lambda$-modules of finite tor-dimension over the small étale site of $X$:
@@
is then a six-functors formalism that satisfies all properties listed in paragraph 1.1. We observe that the dualizable objects are then exactly those complexes whose cohomology objects are what Deligne calls “faisceaux localement constants tordus”. Indeed, the functor
@@
Let $\mathcal{U}$ be an ultrafilter on the set of prime numbers. We define for any noetherian $X$:
@@
(where the colimit and the product are taken in the $\infty$-category of (stable) $\infty$-categories). We define $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})$ as the full subcategory of $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ by noetherian induction: an object $F$ of $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ belongs to $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})$ if there is a dense open subscheme $V$ of $U$ such that the restriction $F_{|_{V}}$ is dualizable in $D^{b}_{c}(V,\mathbf{F}_{\mathcal{V}})^{\prime}$ and $F$ is constructible on a complement of $V$. We define finally
@@
If the set of residue characteristics of $X$ is finite, $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ coincides with the $\mathcal{U}$-indexed ultraproduct of the ordinary categories of constructible sheaves $D^{b}_{c}(X,\mathbf{F}_{\ell})$. If the set of residue characteristics is finite (e.g. is X is of finite type over a semi-local scheme), then $D^{b}_{c}(X,\mathbf{F}_{\mathcal{U}})^{\prime}$ coincides with the ultraproduct of the usual categories of étale sheaves $D(X,\mathbf{F}_{\ell})$, where $\ell$ runs over all primes distinct from the residue characteristics of $X$. The corresponding cohomology is then the ultraproduct in the usual sense of étale cohomology groups:
@@
The only non-trivial property of 1.1 that is difficult to check is (i). In the case of quasi-excellent schemes over a semi-local noetherian ring, this follows right away from Orgogozo’s uniform constructibility theorem [Org19, Théorème 3.1.1].
@@
that are compatible with pullbacks $f^{*}$ for all $f$ as well with direct image with compact support $f_{!}$ for any $f$ separated and of finite presentation. In particular, on noetherian schemes, such morphisms must preserve constructible objets (they preserve dualizable objects, which implies that they preserve constructible objects by noetherian induction). One can prove that $\mathit{DM}$ is the initial object among $\mathbf{Z}$-linear six-functors formalisms satisfying étale descent [^2] One can see that $\mathit{DM}_{h}$ is initial among $\mathbf{Z}$-linear six-functors formalisms satisfying étale hyperdescent: since we know what is the initial six-functors formalism by virtue of Drew and Gallauer’s [DG22, Theorem 7.14], we see easily that it suffices to prove it for $\mathbf{Q}$-linear coefficients, which follows right away from [CD19, Theorem 16.1.2 and 16.2.18], and for finite coefficients, which follows from the combination of rigidity theorems: the one recalled in Example 1.2 together with those stated in [Ayo14, Bac21]. This implies that $\mathit{DM}$ is the initial object among $\mathbf{Z}$-linear six-functors formalisms satisfying the properties of 1.1. For $X$ noetherian with residue fields that have a uniform bound on their Galois cohomological dimension, $\mathit{DM}$ and $\mathit{DM}_{h}$ agree anyway.. In particular, for any other six-functor formalism $D$, there is a unique morphism
@@
We call an object $F$ of $D(X)$ strictly of geometric origin if there is a constructible motivic sheaf $M$ in $\mathit{DM}_{c}(X)$ such that $F\cong r_{D}(M)$ in $D(X)$ (in particular, such an $F$ is constructible). We will say that $D$ behaves well over $S$ if the functor $r_{D}$ commutes with functors of the form $f_{*}$ for any morphism of finite presentation $f$. All examples of six-functors formalisms given in the present notes are known to behave well: in the case where $D$ is $\mathbf{Q}$-linear and defined on schemes of finite type over a field, this is subsumed by [CD19, Theorem 4.4.25] – see Remark 1.13 below to drop the $\mathbf{Q}$-linearity. For $D=D(-,\Lambda)$ with $\Lambda$ of positive characteristic, this is obvious, by rigidity. This is also true in the case where $D(X)=D(X,\mathbf{Z}_{\ell})$ for any $S$ (see [CD16, Theorem 7.2.11]), or $D(X)=D(X,\mathbf{F}_{\mathcal{U}})$ (exercise).
@@
The following statement, is a generalization of Deligne’s generic base change theorem for torsion étale sheaves [Del77, Th. Finitude, 1.9]. With this new degree of generality, it may be seen as an independence of $D$ (in particular, independence of $\ell$) result as well; see Corollary 1.14.
@@
The proof of the generic base change formula really uses étale descent on constructible sheaves: this implies that, for any quasi-finite morphism $f$. the functors $f_{!}$ and $f_{*}$ are conservative, which is a crucial ingredient in Deligne’s proof. It is not known whether the generic base change formula is true or not for more universal theories of motivic sheaves in which we only have Nisnevich descent, such as Voevodsky’s and Morel’s $\mathit{SH}$.
Remark 1.13.
In fact, a close examination of the proof of the generic base change formula shows that, for a given constructible $F$ in $D(X)$ and a morphism of finite presentation $f:X\to Y$ over some base $S$, one can choose the open subset $U\subset S$ so that the conclusion of the generic base change theorem holds for $f_{*}(F)$ in $D$ and so that, for any morphism of six-functors formalisms $\varphi\colon D\to D^{\prime}$ (with $D^{\prime}$ satisfying the assumption of 1.1 as well), the canonical map
$\varphi(f_{*}(F))\to f_{*}(\varphi(F))$
becomes an isomorphism over $U$ and the conclusion of the generic base change theorem holds for $\varphi(F)$ with respect to $f$ in $D^{\prime}$ (the idea is that we can do the proof of the generic base change formula for $F$ in $D$ and for $\varphi(F)$ in $D^{\prime}$ in parallel and observe that $\varphi$ is compatible with each reduction step, thus giving rise to the same choice of locus $U$ over which the base change formula holds).
@@
If moreover $F$ is strictly of geometric origin, then the open subscheme $U\subset S$ in the statement of Theorem 1.11 may be chosen indepently of $D$.
@@
Let $M$ in $\mathit{DM}_{c}(X)$ such that $F\cong r_{D}(M)$ in $D(X)$. Apply the generic base change formula to $M$, which gives an open subscheme $U\subset S$ such that the conclusion of Theorem 1.11 holds in $\mathit{DM}$. The conclusion of Theorem 1.11 holds for $r_{D}(M)$ in $D$ as well, by the previous Remark. [^3] For the reader who does not want to verify the claims of Remark 1.13, Corollary 1.14 can be proved as follows in the case where $D$ behaves well: the conclusion of Theorem 1.11 holds for $F=r_{D}(M)$ in $D$ with respect to the motivically chosen $U\subset S$: since $r_{D}\colon\mathit{DM}\to D$ commutes with the six operations, all the base change isomorphisms involving $M$ in $\mathit{DM}$ will yield analogous base change isomorphisms involving $F$ in $D$.
@@
with $F_{T}$ the pullback of $F$ along the projection $X_{T}\to X$. Moreover, if ever $F$ is strictly of geometric origin, the subscheme $U$ can be chosen independently of $D$.
@@
in which the vertical map on the right hand side is an isomorphism by the computation we just did. Since the middle vertical map is the identity, this proves that the vertical map on the left hand side is an isomorphism. The fact the open subscheme $U\subset\mathbf{P}^{\wedge}$ may be chosen independently of $D$ if ever the sheaf $F$ is strictly of geometric origin follows from Corollary 1.14. ∎